OpenStudy (xxdaboixx):

50 1/2 multiplied by 1/4 equals ?

OpenStudy (mathmale):

Various ways to do this: 1) Mult. 50.5 by 0.25. 2) Convert 50 1/2 to an improper fraction and mult that fraction by 1/4

OpenStudy (mathmale):

\[\frac{ 50\frac{ 1 }{ 2 }*1 }{4 }~by~alones\] is valid.

OpenStudy (xxdaboixx):

It's like 50 1/2 * 1/4=

OpenStudy (mathmale):

Yes, of course. what are you trying to say here? My suggestions apply to that 50 1/2 * 1/4.

OpenStudy (xxdaboixx):

?????

OpenStudy (xxdaboixx):

just a simple question

OpenStudy (mathmale):

Between alones and myself we've given you three different options from which to choose. All will produce the same answer if done properly. Please choose whichever expression looks easiest for you and then perform the multiplication. If you need clarification, ask specific questions.

OpenStudy (xxdaboixx):

ok so I think I will just go with ur's

OpenStudy (xxdaboixx):

thx btw

OpenStudy (mathmale):

Fine. Try it out.

OpenStudy (mathmale):

You're very welcome

OpenStudy (xxdaboixx):

can u check my anwser when I am done?